Solve the following problems.
Suppose COVID-19, a pandemic that spread worldwide has infected 55,000 people on August 2021.
Each day, the number of infected is estimated to be 500. Assuming that the number of daily infections
are constant,
1. What is the monthly growth rate of infection k? Assume time t is measured in months.
2. How many persons will be confirmed positive by the end of December 2021?
3. Assuming that there has a turn of events and majority are vaccinated by September 2021. In
addition, the number of cases dropped with the number of recoveries at a rate of 100 per day.
There have been no additional cases after September. How much time t will elapse for the
cases of COVID-19 to drop to 200?
1
Expert's answer
2021-10-04T01:48:03-0400

1). A=P(1+r/100)n

55000=500(1+r/100)31

110= (1+r/100)31

LN110=31LN (1+r/100)

0.15r=LN (1+R/100)

0.16=r/100

r=16%

2). September

A=500(1+16/100)30

=500(1.16)30

= 42,925

October

A=55000

November

A= 42,925

December

A=55000

Total people to be positive December = 42,925+ 55000+42,925+55000

Ans=250,850